The question is:
A motorboat took 3 hours to make a downstream trip with a current of 6 km/hr. The return trip against the same current took 5hrs. Find the speed of the boat in still water.
I am not sure how to solve resistance problems. If you could just tell me the formula or how to get started, that would be great. thanks.

Answer by Nate(3500)

With Current:
rate = r + 6
time = 3
distance = 3(r + 6)
Against Current:
rate = r - 6
time = 5
distance = 5(r - 6)
3(r + 6) = 5(r - 6)
3r + 18 = 5r - 30
48 = 2r
24km/hr = r